OpenStudy (anonymous):

An object is traveling along a linear path according to the equation s(t) = 4t^3 - 3t^2 + 5 where t is measured in seconds and s(t) measured in meters. How far has the object traveled when its acceleration is zero?

OpenStudy (anonymous):

s''(t) = 24t - 6 sub s''(t)=0: 24t - 6 = 0 24t = 6 t = 6/24 = 0.25 s